OpenStudy (anonymous):

Find the derivative of h(r) = (a*r + b)/m. Assume that a, b and m are constants.

OpenStudy (cggurumanjunath):

a

zepdrix (zepdrix):

\[\Large h(r)\quad=\quad\frac{ar+b}{m} \quad=\quad \frac{a}{m}r+\frac{b}{m}\] \[\Large h'(r)\quad=\quad \frac{a}{m}(1)+0\]

OpenStudy (cggurumanjunath):

sorry it is a/m @zepdrix @alisha

OpenStudy (cggurumanjunath):

@piglet

OpenStudy (cggurumanjunath):

@piglet9

OpenStudy (cggurumanjunath):

@alishabeyer